alis:

The sales of homes in a new development have been increasing. In January, 8 homes were sold, in February, 12 homes were sold. In March, 16 homes were sold. This pattern continued for the remainder of the year. What is the explicit rule that can be used to find the number of homes sold in the nth month of the year?

alis:

these are the answer choices: an=āˆ’4n+4 an=4n+4 an=āˆ’4nāˆ’4 an=4nāˆ’4

umm:

B-an=4n+4

alis:

cool, thx so much! but how did you find that? (for future reference)

umm:

You're welcome. I found it on Quizlet, but the answer that I gave you was on Quizlet so I'm guessing since it was a positive it would still be a positive.

alis:

oh okay thanks again :)

umm:

No problemo